(10) 3. Derive the Backpropagation algorithm (only for the output nodes of a multi-layer perceptron) if the activation function is: (C 、σ, where σ is a parameter to be learned for each neuron? Simplify α v) the equations as much as possible. e
Answer #1

Answer:

For the derivation, lets assume certain conventions:

1. Subscript k will represent an output node. Subscript j will represent a hidden node.

2. Weight from a hidden node to an output node will be represented as w_{hj}^{k}.

3. Weight change rule for hidden to output layer weights will be given by:

w_{hj}^{k+1}=w_{hj}^{k} + \Delta w_{hj}^{k}

where \Delta w_{hj}^{k} is the weight change.

4. Activation function or signal will be represented as \varphi (\nu). In present case, \varphi (\nu) = e^{-(\frac{\nu}{\sigma})^{2}} .

5. y_{j}^{k} will be activation value at any output node.

Derivation of backpropagation algorithm for output layer:

Computation of neoronal signals:

A. For output layer, neuronal signal and activation value at any output node can be calculated as:

1. y_{j}^{k} = \sum_{h=0}^{q}w_{hj}^{k}\varphi (\nu_{h}^{k}), j = 1,............,p

2. \varphi (\nu_{h}^{k}) = e^{-(\frac{\nu_{h}^{k}}{\sigma })^{2}}, j = 1,............,p

where w_{0j}^{k} will be the biases of output neurons.

Computation of error gradients:

B. Weight gradients of hidden weight gradients:

It can be calculated using chain rule of calculus as shown below:

3. \frac{\partial \varepsilon _{k}}{\partial w_{hj}^{k}} = \frac{\partial \varepsilon _{k}}{\partial \varphi(y_{j}^{k})}\frac{\partial \varphi(y_{j}^{k})}{\partial y_{j}^{k}}\frac{\partial y_{j}^{k}}{\partial w_{hj}^{k}}

Individual derivatives in the above equation can be calculated as:

i) \frac{\partial \varepsilon _{k}}{\partial \varphi(y_{j}^{k})} = -(d_{j}^{k} - \varphi(y_{j}^{k})) = -e_{j}^{k}

where d_{j}^{k} is the desired output for given node.

ii) \frac{\partial \varphi(y_{j}^{k})}{\partial y_{j}^{k}} = \varphi(y_{j}^{k})(1-\varphi(y_{j}^{k}))

iii) \frac{\partial y_{j}^{k}}{\partial w_{hj}^{k}} = \varphi(z_{h}^{j})

where z_{h}^{j} is activaion at concerned hidden unit.

Now on replacing values from i), ii) and iii) in 3, we get:

\frac{\partial \varepsilon _{k}}{\partial w_{hj}^{k}} = -e_{j}^{k}\varphi(y_{j}^{k})(1-\varphi(y_{j}^{k}))\varphi (z_{h}^{k})

Weight updates for output layer:          

C. For hidden to output layer weights:

w_{hj}^{k+1}=w_{hj}^{k} + \Delta w_{hj}^{k}

           =w_{hj}^{k} + \sigma \left ( -\frac{\partial \varepsilon _{k}}{\partial w_{hj}^{k}} \right )

           = w_{hj}^{k} + \sigma e_{j}^{k}\varphi(y_{j}^{k})(1-\varphi(y_{j}^{k}))\varphi (z_{h}^{k})